### v1

@@ Abstract
Two decades of research on the euro’s trade effects have produced estimates ranging from 4% to 30%, with no consensus on the magnitude. We find evidence that this divergence may reflect genuine heterogeneity in the euro’s trade effect across country pairs rather than methodological differences alone. Using Eurostat data on 15 EU countries from 1995–2015, we estimate that euro adoption increased bilateral trade by 24% on average (15.0% after fixed effects correction), but effects range from $12\%$ to $68\%$ across eurozone pairs. Core eurozone pairs (e.g., Germany-France, Germany-Netherlands) show large gains, while peripheral pairs involving Finland, Greece, and Portugal saw smaller or negative effects, with some negative estimates statistically significant and interpretable as trade diversion. Pre-euro trade intensity and GDP explain over 90% of this variation. Extending to EU28, we find evidence that crisis-era adopters (Slovakia, Estonia, Latvia) pull down naive estimates to 5%, but accounting for fixed effects recovers estimates of 14.0%, consistent with the EU15 fixed-effects baseline of 15.0%. Illustrative counterfactual analysis suggests non-eurozone members would have experienced varied effects: UK (+24%), Sweden (+20%), Denmark (+19%). The wide range of prior estimates appears to be largely a feature of the data, not a bug in the methods.
@@ 3.4 Causal Forest with Double Machine Learning
This approach removes regularization bias from the ML first stage and provides valid confidence intervals. It is doubly robust: consistent if either $\hat{m}$ or $\hat{e}$ is correctly specified.
@@ 3.6 Treatment Definition and Identification
The sample is restricted to pairs where both countries are EU members, ensuring all units faced a credible possibility of treatment. Countries outside the EU (e.g., United States, Japan, Switzerland) could never adopt the euro and thus do not constitute a valid counterfactual. Within the EU sample, eurozone members (Austria, Belgium-Luxembourg, Finland, France, Germany, Greece, Ireland, Italy, Netherlands, Portugal, Spain) constitute the treatment group, while EU members that opted out (Denmark, Sweden, United Kingdom) constitute the control group.
@@
For PPML specifications in the gravity benchmark, we use trade in levels following Santos Silva and Tenreyro (2006). Effect modifiers $X$ — the variables that may drive heterogeneity in treatment effects — include log GDP product, log GDP per capita, and pre-euro trade intensity (average bilateral trade 1990–1998).
@@ 3.7 Threats to Identification
We provide two forms of evidence supporting the parallel trends assumption. First, an event study analysis (Figure A.1) shows that pre-treatment coefficients (1995–1997) are small in magnitude ($3\%$ to $5\%$) compared to post-treatment effects ($8\%$ to $24\%$), with a clear break at 1999 when effects become positive. Second, placebo tests assigning fake treatment dates (1995, 1997) find no significant “effects,” suggesting our estimates do not reflect pre-existing differential trends.
Following Oster (2019), we assess how much selection on unobservables would be required to explain away our results. The analysis suggests that unobserved confounders would need to be substantially more important than the observed confounders (GDP, GDP per capita, year) to reduce the estimated effect to zero. Given that our controls capture the main economic determinants of both euro adoption and trade, this degree of omitted variable bias appears implausible.
Leave-one-out analysis dropping each country in turn shows that no single country drives the results. The ATE remains stable within the confidence interval of the full-sample estimate regardless of which country is excluded, including Luxembourg (which has the largest estimated effects) and peripheral economies like Greece and Portugal.
@@ 4 Data
We use bilateral trade data from Eurostat, which provides comprehensive trade statistics for EU member states. The dataset covers 15 EU countries (EU15) from 1995 to 2015, matching the methodology of Gunnella et al. (2021). We restrict the sample to the EU15 (pre-2004 enlargement members) for two reasons. First, recent euro adopters among the 2004+ enlargement countries have insufficient post-treatment data: Slovenia joined in 2007 (8 years), Slovakia in 2009 (6 years), Estonia in 2011 (4 years), Latvia in 2014 (1 year), and Lithuania in 2015 (0 years in our sample). Second, this sample matches the sample used by Gunnella et al. (2021), enabling direct comparison. The outcome, treatment, and control variables are defined in Section 3.4.
@@
Table 2 reports summary statistics. Log real trade has a mean of 22.31 with substantial variation (standard deviation of 1.73), reflecting heterogeneity in bilateral trade flows across EU pairs. The euro adoption indicator equals one for 50% of observations, reflecting the balanced panel structure with eurozone pairs observed both before and after 1999. Table 3 shows the panel contains 2,189 observations covering 105 unique country pairs across 15 countries. Of these, 1,100 (50.3%) are treated while 1,089 (49.7%) serve as controls.
@@
Log GDP per capita 20.48 0.65 18.31 22.45
EU membership 0.99 0.08 0.00 1.00
Euro adoption 0.50 0.50 0.00 1.00
@@
Year range 1988–2015
Treated (euro = 1) 1,100 (50.3%)
Control (euro = 0) 1,089 (49.7%)
@@
Pre-euro trade intensity 22.10 22.20 -0.10 -0.06
Note: Standardized difference >> 0.25 indicates meaningful imbalance. Eurozone = pairs where both countries adopted the euro; Non-EZ = pairs involving Denmark, Sweden, or UK. Pre-euro trade intensity is the average log bilateral trade during 1990–1998. Comparison uses post-1999 observations only.
@@
To further assess overlap between treated and control groups, we estimate propensity scores using logistic regression with log GDP product, log GDP per capita, and pre-euro trade intensity as predictors. Figure 5 shows the distribution of propensity scores by treatment status. Both treated (eurozone) and control (non-eurozone) pairs span a wide range of propensity scores with substantial overlap, supporting the positivity assumption required for causal inference. Table 5 provides summary statistics.
@@
N observations 1,159 792
Mean propensity score 0.624 0.550
Std. dev. 0.128 0.129
Min 0.278 0.252
Max 0.931 0.875
Common support region: [0.278, 0.875]
In common support 1,143 (98.6%) 787 (99.4%)
@@
The CausalForestDML is implemented with Random Forest first-stage models (200 trees, min_samples_leaf = 20), a causal forest with 500 trees and honest splitting (min_samples_leaf = 30), and a Random Forest Classifier for the binary treatment model.
@@ 5.1 Gravity Benchmark
Before examining heterogeneity, we establish a gravity benchmark using the canonical specification from the euro trade literature. Table 6 presents estimates from several specifications on the same EU15 sample. The two-way fixed effects OLS estimate is 16.3%, while PPML gravity with pair and year fixed effects yields 9.9%. These estimates fall within the 4–30% range documented in prior literature (Rose, 2016; Gunnella et al., 2021).
Table 6: Comparison of Euro Trade Effect Estimates (EU15 Sample)
@@
Two-way FE OLS Pair + Year FE 0.151 +16.3% [0.046, 0.257]
PPML Gravity Pair + Year FE 0.094 +9.9% [-0.000, 0.188]
PPML Gravity Exp + Imp + Year FE 0.338 +40.3% [0.032, 0.645]
Causal Forest (DML) Heterogeneous effects 0.219 +24.5% [0.063, 0.375]
CFFE Node-level FE 0.141 +15.1% [0.103, 0.178]
Notes: All models estimated on EU15 bilateral trade data, 1995-2015. Sample includes 15 EU member states with 3 non-eurozone controls (Denmark, Sweden, UK). PPML = Poisson Pseudo-Maximum Likelihood following Santos Silva & Tenreyro (2006). DML = Double Machine Learning. CFFE = Causal Forests with Fixed Effects. Standard errors clustered at the country-pair level.
The causal forest DML estimate of 24% is somewhat higher than the gravity estimates. This difference reflects two factors. First, the causal forest uses flexible nonparametric first-stage estimation rather than linear fixed effects, potentially capturing nonlinear relationships between confounders and outcomes. Second, and more importantly, the causal forest estimates a different object: while gravity models estimate a single average effect constrained to be constant across pairs, the causal forest estimates heterogeneous effects that can vary with pair characteristics. The ATE from the causal forest is a weighted average of these heterogeneous effects, where the weights depend on the covariate distribution.
The CFFE estimate of 15.0% falls between the gravity and naive causal forest estimates, suggesting that proper fixed effects handling moderates the causal forest estimate toward the gravity benchmark. This convergence is reassuring: when we account for pair and year fixed effects within the causal forest framework, we recover estimates consistent with the gravity literature.
@@ 5.3 Heterogeneous Treatment Effects
Euro adoption increased bilateral trade by 24% on average (95% CI: $6\%,46\%]$) using the naive causal forest without fixed effects. [^2] Since the outcome variable is log trade, the coefficient estimate is in log points. We convert to percentage change using $(\exp(\hat{\tau})-1)\times 100$. For example, an estimate of 0.219 log points corresponds to $(\exp(0.219)-1)\times 100=24\%$. Figures report estimates in log points; text reports percentage changes. This estimate falls between the gravity model range (4–6%) and the SCM estimates ($\sim$30%) reported by Gunnella et al. (2021). The confidence interval is tighter than in previous studies, reflecting the larger sample and more precise estimation.
The CATE distribution reveals heterogeneity that cannot be captured by methods estimating a single average effect. Among eurozone pairs, effects range from $12\%$ to $68\%$, with substantial variation across pairs. The euro does not have a single effect on trade. Different country pairs experienced dramatically different impacts, from near-zero to substantial gains. Figure 6 shows the distribution for eurozone pairs.
Figure 6: Distribution of Conditional Average Treatment Effects (CATE) for eurozone country pairs. The red dashed line indicates the Average Treatment Effect (ATE). Effects range from $12\%$ to $68\%$ across pairs.
@@
Luxembourg $\leftrightarrow$ Netherlands +68.2
France $\leftrightarrow$ Luxembourg +68.0
Germany $\leftrightarrow$ Luxembourg +67.3
Italy $\leftrightarrow$ Luxembourg +65.5
Belgium $\leftrightarrow$ Luxembourg +62.1
@@
Greece $\leftrightarrow$ Ireland -9.8
Ireland $\leftrightarrow$ Portugal -10.6
Finland $\leftrightarrow$ Greece -11.1
Greece $\leftrightarrow$ Portugal -12.2
Finland $\leftrightarrow$ Portugal -12.2
Table 7 shows the top and bottom eurozone pairs by estimated euro effect. The highest effects are concentrated among pairs involving Luxembourg (Luxembourg–Netherlands at +68%, France–Luxembourg at +68%, Germany–Luxembourg at +67%, Italy–Luxembourg at +66%). The lowest effects — including some near-zero point estimates — involve peripheral pairs such as Finland–Portugal and Greece–Portugal. Table 8 provides confidence intervals for all pair-level effects. Notably, while some point estimates are negative, none are statistically distinguishable from zero at the 5% level, suggesting these pairs experienced no detectable euro effect rather than genuinely negative effects.
@@
Luxembourg $\leftrightarrow$ Netherlands +68.2% [+27.9%, +121.0%] Yes
France $\leftrightarrow$ Luxembourg +68.0% [+37.8%, +104.8%] Yes
Germany $\leftrightarrow$ Luxembourg +67.3% [+39.3%, +100.9%] Yes
Italy $\leftrightarrow$ Luxembourg +65.5% [+34.0%, +104.4%] Yes
Belgium $\leftrightarrow$ Luxembourg +62.1% [+18.3%, +122.0%] Yes
Ireland $\leftrightarrow$ Luxembourg +60.6% [+12.8%, +128.6%] Yes
Austria $\leftrightarrow$ Luxembourg +60.6% [+15.8%, +122.7%] Yes
Luxembourg $\leftrightarrow$ Spain +59.0% [+26.5%, +100.0%] Yes
Finland $\leftrightarrow$ Luxembourg +57.2% [+10.6%, +123.6%] Yes
Belgium $\leftrightarrow$ France +55.2% [+43.0%, +68.5%] Yes
@@
Austria $\leftrightarrow$ Greece +8.7% [-0.0%, +18.2%] No
Austria $\leftrightarrow$ Finland +3.5% [-8.0%, +16.4%] No
Austria $\leftrightarrow$ Ireland +2.8% [-7.3%, +14.1%] No
Austria $\leftrightarrow$ Portugal +0.3% [-7.8%, +9.2%] No
Finland $\leftrightarrow$ Ireland -10.2% [-20.1%, +0.8%] No
Greece $\leftrightarrow$ Ireland -11.8% [-20.1%, -2.6%] Yes (-)
Finland $\leftrightarrow$ Greece -11.9% [-19.9%, -3.2%] Yes (-)
Ireland $\leftrightarrow$ Portugal -12.4% [-21.0%, -2.9%] Yes (-)
Greece $\leftrightarrow$ Portugal -13.8% [-21.7%, -5.1%] Yes (-)
Finland $\leftrightarrow$ Portugal -14.0% [-22.4%, -4.7%] Yes (-)
@@
Table 9 aggregates the pair-level effects to show the average euro effect for each eurozone member country. Luxembourg (+60%), Belgium (+40%), and Netherlands (+39%) show the largest average effects, while Greece (+11%) and Portugal (+13%) show the smallest. This core-periphery pattern suggests the euro’s benefits were not uniformly distributed.
@@
Luxembourg +59.5 0.09 +40.1 +82.0
Belgium +39.6 0.12 -9.1 +78.3
Netherlands +39.2 0.13 +12.6 +78.4
Germany +38.2 0.15 +9.1 +82.0
France +30.8 0.15 +12.6 +81.8
Italy +28.3 0.14 +13.2 +81.7
Austria +22.6 0.18 -15.5 +77.9
Spain +22.1 0.10 +10.0 +74.7
Ireland +17.5 0.22 -15.5 +77.7
Finland +14.5 0.21 -15.2 +66.2
Portugal +13.2 0.20 -15.8 +48.5
Greece +11.1 0.17 -15.8 +48.8

### v2

@@ Abstract
Two decades of research on the euro’s trade effects have produced estimates ranging from 4% to 30%, with no consensus on the magnitude. We find evidence that this divergence may reflect genuine heterogeneity in the euro’s trade effect across country pairs rather than methodological differences alone. Using Eurostat data on 15 EU countries (12 eurozone members plus Denmark, Sweden, and the UK as controls) from 1995–2015, we estimate that euro adoption increased bilateral trade by 29% on average (14.1% after fixed effects correction), but effects range from $12\%$ to $79\%$ across eurozone pairs. Core eurozone pairs (e.g., Germany–France, Germany–Netherlands) show large gains, while peripheral pairs involving Finland, Greece, and Portugal saw smaller or negative effects, with some negative estimates statistically significant and interpretable as trade diversion. Pre-euro trade intensity and GDP account for over 90% of feature importance in explaining this heterogeneity. Extending to EU28, we find evidence that crisis-era adopters (Slovakia, Estonia, Latvia) pull down naive estimates to 4.3%, but accounting for fixed effects recovers estimates of 13.4%, consistent with the EU15 fixed-effects baseline of 14.1%. Illustrative counterfactual analysis suggests non-eurozone members would have experienced varied effects: UK (+33%), Sweden (+22%), Denmark (+19%). The wide range of prior estimates appears to be largely a feature of the data, not a bug in the methods.
@@ 3.4 Causal Forest with Double Machine Learning
This approach removes regularization bias from the ML first stage and provides valid confidence intervals. It is doubly robust: consistent if either $\hat{m}$ or $\hat{e}$ is correctly specified. We implement CausalForestDML with Random Forest first-stage models (200 trees, min_samples_leaf = 20), a causal forest with 500 trees and honest splitting (min_samples_leaf = 30), and a Random Forest Classifier for the binary treatment model.
@@ 3.6 Treatment Definition and Identification
The sample is restricted to pairs where both countries are EU members, ensuring all units faced a credible possibility of treatment. Countries outside the EU (e.g., United States, Japan, Switzerland) could never adopt the euro and thus do not constitute a valid counterfactual. Within the EU sample, eurozone members (Austria, Belgium, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Portugal, Spain) constitute the treatment group, while EU members that opted out (Denmark, Sweden, United Kingdom) constitute the control group.
@@
For PPML specifications in the gravity benchmark, we use trade in levels following Santos Silva and Tenreyro (2006). Effect modifiers $X$ — the variables that may drive heterogeneity in treatment effects — include log GDP product, log GDP per capita, and pre-euro trade intensity (average bilateral trade 1995–1998, the pre-treatment period in our sample).
@@ 3.7 Threats to Identification
We provide two forms of evidence supporting the parallel trends assumption. First, an event study analysis (Figure A.1) shows that pre-treatment coefficients (1995–1997) are small in magnitude ($3\%$ to $5\%$) compared to post-treatment effects ($8\%$ to $24\%$), with a clear break at 1999 when effects become positive. Second, placebo tests (Figure A.2) assigning fake treatment dates (1995, 1997) find no significant “effects,” suggesting our estimates do not reflect pre-existing differential trends.
Following Oster (2019), we assess how much selection on unobservables would be required to explain away our results (Table A.3). The analysis suggests that unobserved confounders would need to be substantially more important than the observed confounders (GDP, GDP per capita, year) to reduce the estimated effect to zero. Given that our controls capture the main economic determinants of both euro adoption and trade, this degree of omitted variable bias appears implausible.
Leave-one-out analysis (Figure A.4) dropping each country in turn shows that no single country drives the results. The ATE remains stable within the confidence interval of the full-sample estimate regardless of which country is excluded, including Luxembourg (which has the largest estimated effects) and peripheral economies like Greece and Portugal.
@@ 4 Data
We use bilateral trade data from Eurostat, which provides comprehensive trade statistics for EU member states. The dataset covers 15 EU countries (EU15) from 1995 to 2015, matching the methodology of Gunnella et al. (2021). While Eurostat trade data extends back to 1988, GDP data from Eurostat’s national accounts becomes consistently available only from 1995, making earlier years unsuitable for our analysis which requires GDP controls. We restrict the sample to the EU15 (pre-2004 enlargement members) for two reasons. First, recent euro adopters among the 2004+ enlargement countries have insufficient post-treatment data: Slovenia joined in 2007 (8 years), Slovakia in 2009 (6 years), Estonia in 2011 (4 years), Latvia in 2014 (1 year), and Lithuania in 2015 (0 years in our sample). Second, this sample matches the sample used by Gunnella et al. (2021), enabling direct comparison. The outcome, treatment, and control variables are defined in Section 3.4.
@@
Table 2 reports summary statistics. Log real trade has a mean of 22.30 with substantial variation (standard deviation of 1.73), reflecting heterogeneity in bilateral trade flows across EU pairs. The euro adoption indicator equals one for 51% of observations, reflecting the balanced panel structure with eurozone pairs observed both before and after 1999. Table 3 shows the panel contains 2,149 observations covering 105 unique country pairs across 15 countries. Of these, 1,100 (51.2%) are treated while 1,049 (48.8%) serve as controls.
@@
Log GDP per capita 20.50 0.64 18.31 22.45
EU membership 1.00 0.00 1.00 1.00
Euro adoption 0.51 0.50 0.00 1.00
@@
Year range 1995–2015
Treated (euro = 1) 1,100 (51.2%)
Control (euro = 0) 1,049 (48.8%)
@@
Pre-euro trade intensity 22.10 22.23 -0.13 -0.09
Note: Standardized difference >> 0.25 indicates meaningful imbalance. Eurozone = pairs where both countries adopted the euro; Non-EZ = pairs involving Denmark, Sweden, or UK. Pre-euro trade intensity is the average log bilateral trade during 1995–1998. Comparison uses post-1999 observations only.
@@
To further assess overlap between treated and control groups, we estimate propensity scores using both logistic regression and random forest with log GDP product, log GDP per capita, and pre-euro trade intensity as predictors. Figure 5 shows the distribution of propensity scores by treatment status. The logistic regression estimates (left panel) show substantial overlap, while the random forest estimates (right panel) show tighter separation—reflecting RF’s ability to capture nonlinear relationships in treatment assignment. Both treated (eurozone) and control (non-eurozone) pairs span a wide range of propensity scores under both methods, supporting the positivity assumption required for causal inference. The tighter RF distribution is relevant because our DML approach uses flexible ML models in the first stage; the presence of overlap even under RF’s more precise fit is reassuring. Table 5 provides summary statistics.
@@
N observations 1,155 756
Mean propensity score 0.640 0.550
Std. dev. 0.140 0.139
Min 0.256 0.223
Max 0.950 0.904
Common support region: [0.256, 0.904]
In common support 1,139 (98.6%) 753 (99.6%)
@@ 5.1 Gravity Benchmark
Before examining heterogeneity, we establish a gravity benchmark using the canonical specification from the euro trade literature. Table 6 presents estimates from several specifications on the same EU15 sample. The two-way fixed effects OLS estimate is 17.0%, while PPML gravity with pair and year fixed effects yields 12.8%. The three-way fixed effects PPML specification—with exporter-time, importer-time, and pair fixed effects following structural gravity best practices (Anderson and van Wincoop, 2003)—yields a higher estimate of 22.6%. These estimates fall within the 4–30% range documented in prior literature (Rose, 2016; Gunnella et al., 2021).
Table 6: Comparison of Euro Trade Effect Estimates
@@
Two-way FE OLS Pair + Year FE 0.157 +17.0% [0.048, 0.267]
PPML Gravity Pair + Year FE 0.120 +12.8% [0.035, 0.205]
PPML Gravity Three-way FE 0.204 +22.6% [0.123, 0.285]
Causal Forest DML 0.252 +28.6% [0.103, 0.401]
CFFE Node-level FE 0.133 +14.2% [0.103, 0.162]
Notes: All models estimated on EU15 bilateral trade data, 1995–2015. Two-way FE OLS uses pair and year fixed effects with clustered standard errors. PPML = Poisson Pseudo-Maximum Likelihood following Santos Silva & Tenreyro (2006). Three-way FE includes exporter-time, importer-time, and pair fixed effects (Head & Mayer, 2014). Causal Forest uses Double Machine Learning for heterogeneous treatment effects. CFFE = Causal Forests with Fixed Effects. Standard errors clustered at the country-pair level.
The causal forest DML estimate of 29% is somewhat higher than the gravity estimates. This difference reflects two factors. First, the causal forest uses flexible nonparametric first-stage estimation rather than linear fixed effects, potentially capturing nonlinear relationships between confounders and outcomes. Second, and more importantly, the causal forest estimates a different object: while gravity models estimate a single average effect constrained to be constant across pairs, the causal forest estimates heterogeneous effects that can vary with pair characteristics. The ATE from the causal forest is a weighted average of these heterogeneous effects, where the weights depend on the covariate distribution.
The CFFE estimate of 14.1% falls between the gravity and naive causal forest estimates, suggesting that proper fixed effects handling moderates the causal forest estimate toward the gravity benchmark. This convergence is reassuring: when we account for pair and year fixed effects within the causal forest framework, we recover estimates consistent with the gravity literature.
@@ 5.3 Heterogeneous Treatment Effects
Euro adoption increased bilateral trade by 29% on average (95% CI: $11\%,49\%]$) using the naive causal forest without fixed effects. [^2] Since the outcome variable is log trade, the coefficient estimate is in log points. We convert to percentage change using $(\exp(\hat{\tau})-1)\times 100$. For example, an estimate of 0.252 log points corresponds to $(\exp(0.252)-1)\times 100=29\%$. Figures report estimates in log points; text reports percentage changes. This estimate falls between the gravity model range (4–6%) and the SCM estimates ($\sim$30%) reported by Gunnella et al. (2021). The confidence interval is tighter than in previous studies, reflecting the larger sample and more precise estimation.
The CATE distribution reveals heterogeneity that cannot be captured by methods estimating a single average effect. Among eurozone pairs, effects range from $12\%$ to $79\%$, with substantial variation across pairs. The euro does not have a single effect on trade. Different country pairs experienced dramatically different impacts, from near-zero to substantial gains. Figure 6 shows the distribution for eurozone pairs.
Figure 6: Distribution of Conditional Average Treatment Effects (CATE) for eurozone country pairs. The red dashed line indicates the Average Treatment Effect (ATE). Effects range from $12\%$ to $79\%$ across pairs.
@@
France $\leftrightarrow$ Italy +78.5
Germany $\leftrightarrow$ Italy +78.5
Belgium $\leftrightarrow$ Germany +78.5
Austria $\leftrightarrow$ Germany +78.4
France $\leftrightarrow$ Germany +78.4
@@
Greece $\leftrightarrow$ Ireland -9.6
Ireland $\leftrightarrow$ Portugal -10.4
Finland $\leftrightarrow$ Greece -11.4
Greece $\leftrightarrow$ Portugal -12.0
Finland $\leftrightarrow$ Portugal -12.3
Table 7 shows the top and bottom eurozone pairs by estimated euro effect. The highest effects are concentrated among core European pairs (France–Italy at +79%, Germany–Italy at +78%, Belgium–Germany at +78%, Austria–Germany at +78%, France–Germany at +78%). The lowest effects — including some negative point estimates — involve peripheral pairs such as Greece–Ireland ($10\%$) and Ireland–Portugal ($10\%$). Table 8 provides confidence intervals for all pair-level effects. Notably, while some point estimates are negative, most are not statistically distinguishable from zero at the 5% level. However, two pairs—Greece–Portugal and Finland–Portugal—show statistically significant negative effects, suggesting genuine trade diversion for these peripheral pairs.
@@
France $\leftrightarrow$ Italy +78.5% [+60.8%, +98.2%] Yes
Germany $\leftrightarrow$ Italy +78.5% [+60.7%, +98.2%] Yes
Belgium $\leftrightarrow$ France +78.5% [+60.2%, +98.9%] Yes
Belgium $\leftrightarrow$ Germany +78.4% [+60.3%, +98.5%] Yes
Austria $\leftrightarrow$ Germany +78.4% [+60.2%, +98.6%] Yes
France $\leftrightarrow$ Germany +78.4% [+60.6%, +98.1%] Yes
Germany $\leftrightarrow$ Netherlands +78.2% [+60.5%, +97.8%] Yes
Ireland $\leftrightarrow$ Luxembourg +77.5% [+34.2%, +134.6%] Yes
Austria $\leftrightarrow$ Luxembourg +77.1% [+32.4%, +137.0%] Yes
Belgium $\leftrightarrow$ Luxembourg +76.6% [+30.7%, +138.6%] Yes
@@
Austria $\leftrightarrow$ Greece -2.2% [-11.0%, +7.5%] No
Austria $\leftrightarrow$ Finland -3.4% [-13.2%, +7.6%] No
Austria $\leftrightarrow$ Ireland -4.6% [-15.4%, +7.7%] No
Austria $\leftrightarrow$ Portugal -6.4% [-15.6%, +3.8%] No
Finland $\leftrightarrow$ Ireland -9.9% [-22.9%, +5.3%] No
Ireland $\leftrightarrow$ Portugal -13.4% [-26.7%, +2.3%] No
Greece $\leftrightarrow$ Ireland -13.4% [-25.8%, +1.0%] No
Finland $\leftrightarrow$ Greece -13.6% [-25.5%, +0.1%] No
Greece $\leftrightarrow$ Portugal -15.1% [-26.6%, -1.9%] Yes (-)
Finland $\leftrightarrow$ Portugal -15.2% [-27.9%, -0.3%] Yes (-)
@@
Table 9 aggregates the pair-level effects to show the average euro effect for each eurozone member country. Luxembourg (+69%), Germany (+50%), Belgium (+47%), and Netherlands (+47%) show the largest average effects, while Greece (+10%) and Finland (+15%) show the smallest. This core-periphery pattern suggests the euro’s benefits were not uniformly distributed.
@@
Luxembourg +69.0 0.06 +42.7 +85.9
Germany +49.5 0.19 +11.8 +80.3
Belgium +47.3 0.16 -4.9 +85.9
Netherlands +47.3 0.13 +18.9 +82.4
France +36.7 0.20 +8.1 +80.2
Italy +34.8 0.16 +13.2 +78.6
Spain +26.0 0.10 +8.8 +73.3
Austria +22.5 0.24 -17.1 +85.7
Ireland +18.8 0.25 -17.3 +85.7
Portugal +15.8 0.23 -17.4 +71.1
Finland +14.6 0.23 -17.5 +80.3
Greece +9.6 0.21 -17.5 +71.1